### v1

# Efficient Transfer Learning for Video-language Foundation Models

Abstract
[…]
1 Introduction
Figure 1: We compared the number of trainable parameters of our method with other state-of-the-art methods, as well as the average Harmonic Mean Performance on the base-to-novel generalization task across four datasets. It can be observed that our method uses fewer parameters while achieving state-of-the-art results.
Multi-modal foundation models (e.g., CLIP [28], ViCLIP [36], BLIP2 [21]) that are trained on large-scale datasets of paired multi-modal data have shown exceptional generalization capabilities. The semantic visual concepts empowered by large-scale multi-modal alignment can be widely applied to various downstream tasks, such as few-shot image/video classification [19, 3, 43, 29], open-vocabulary detection [12, 39], and segmentation [23]. In this paper, we focus on transferring the pre-trained multi-modal foundation models model (i.e., CLIP [28] and ViCLIP [36]) to the domain of video action recognition, emphasizing its crucial role in driving advancements in this field.
The key to the adaptation process of foundational multi-modal models lies in injecting domain-specific expertise. For action recognition tasks, this need is reflected in the dynamic nature of video data, which requires models to effectively understand context and temporal relationships. Consequently, existing methods primarily leverage pre-trained CLIP models, enhancing them with carefully designed prompts [37], adapters [44, 25, 41], and temporal modules [24] to enable the model to acquire more domain-specific knowledge about videos. For example, XCLIP [24] introduces cross-frame attention and multi-frame integration modules for temporal modeling. ActionCLIP [34], by utilizing pre-training to obtain powerful representations, applies prompt engineering to bring action recognition tasks closer to the pre-trained model, followed by end-to-end fine-tuning on the target dataset to achieve superior performance. Although the model capacity brought by a large number of learnable parameters enables better fitting of video-specific inductive biases, it may catastrophically forget the original generalization knowledge, thus reducing the generalizability of models.
On the other hand, to enhance the performance of CLIP on video understanding tasks, Wang et al. [36] proposed ViCLIP, which builds upon CLIP and is trained on a large-scale video-text pair dataset. ViCLIP replaces the native attention mechanism in the visual encoder with spatiotemporal attention, while keeping other design components unchanged. However, to date, no work has explored how to leverage ViCLIP for efficient transfer learning, and methods designed for CLIP cannot be directly applied to ViCLIP.
Thus, is there a method that can efficiently fine-tune pre-trained model without damaging its generalizability? A natural idea is to directly apply the efficient transfer methods of image/text foundation models (i.e., LoRA [17] and AdaptFormer [5]) to video tasks. However, these methods are all based on uni-modal information, making it difficult to adapt to multi-modal models. A simple approach is to apply adapters separately to the two modalities. However, this method overlooks the relationship between text and video representations. Therefore, directly applying the same adapters may not be sufficient to capture task-specific cues that vary between vision and language. Furthermore, existing methods do not fully consider the characteristics of text and video representations. Transfer learning often involves a trade-off between discrimination and generalization—specifically, fine-tuning features that are task-specific and discriminative, while preserving those that are more broadly applicable across tasks.
To this end, we propose a novel Mutli-modal Spatio-Temporal Adapter (MSTA) architecture for multi-modal foundation models such that the text and video representations can be better aligned. As shown in Table 1, our method achieves the best generalization performance with a minimal number of parameters. Specifically, our MSTA includes independent projection layers for the text and video branches to learn task-specific knowledge for each modality. To achieve effective alignment between modalities, we introduce a shared unified feature projection layer, which is jointly utilized by both modalities. During the fine-tuning phase, this shared feature space can receive gradients from both modalities [40], thereby optimizing the alignment between them. For the video branch, we further design two simple up-projection layers to enhance the adaptation capabilities for temporal and spatial features. To further mitigate over-fitting and enhance the generalizability, we introduce a spatio-temporal description-guided consistency constraint. This method successfully transfers knowledge from a frozen encoder to a learnable encoder through knowledge distillation, allowing the model to maintain the generalization strength of the pre-trained base model when handling new tasks in few-shot scenarios [30]. Specifically, we impose a consistency constraint on the text branch between the trainable model with MSTA and the pre-trained model. We leverage a pre-trained large language model (e.g., DeepSeek [7]) to generate more detailed and descriptive sentences of temporal and spatial features, and impose a consistency constraint between the representations of the learnable text encoder and the pre-trained text encoder based on these sentences. To validate the effectiveness of our model, we perform extensive experiments across six benchmark datasets: Kinetics-400 [1], Kinetics-600 [2], UCF-101 [31], HMDB-51 [20], SomethingSomething V2 [13], and ActivityNet [15]. The results show that our approach attains state-of-the-art performance in open-vocabulary tasks, such as zero-shot and few-shot learning, and consistently enhances performance when integrated with existing pipelines in fully supervised settings. The key contributions of this work are as follows:
  • We propose a novel Multi-modal Spatio-Temporal Adapter (MSTA) that contains separate projection layers to improve feature representations for video and text encoders independently. Additionally, we implement a shared projection layer to improve the alignment between video and language representations.
  • We propose a spatio-temporal description-guided consistency constraint for large multi-modal foundation models, enabling them to learn new tasks from a small number of samples while maintaining their generalization capability.
  • Extensive experiments demonstrate that MSTA achieves an optimal balance between new and old knowledge while training only a small number of parameters. Comprehensive ablation studies showcase the scalability and effectiveness of our proposed method.
2 Related Works
Multi-modal Foundation Models. The latest advances in multi-modal foundation models have significantly impacted the field of computer vision, especially in tasks that combine language and vision. Representative models include, but are not limited to, CLIP [28], BLIP [21], Florence [42], Kosmos [26], InternVideo [35], and ViCLIP [36]. These models leverage self-supervised paradigms extracted from large-scale multi-modal web data for training. For example, CLIP [28] is trained using contrastive loss [33] on approximately 400 million image-text pairs, while ViCLIP [36] is trained on about 10 million video-text pairs. By collecting more multi-modal data, these models have demonstrated promising performance across various downstream applications. Despite their ability to learn generalized representations, effectively adapting these pre-trained models to specific downstream tasks remains a major challenge, particularly in few-shot settings. To address this, many studies have proposed various methods tailored to different tasks, such as few-shot action recognition [34, 29], video question answering [11], and segmentation [9]. In contrast, this work proposes a novel multi-modal spatio-temporal adapter to effectively adapt multi-modal foundation models for generalization tasks.
Efficient Transfer Learning. Traditional approaches [8, 10] migrate models to downstream tasks by fine-tuning all parameters of the pre-trained network. However, as model sizes increase, this traditional paradigm faces a significant computational burden, and fine-tuning a large number of parameters often leads to severe overfitting, especially in low-sample scenarios. Recently, numerous methods [19, 3, 43, 29] have been proposed to explore the adaptation of pre-trained vision-language models [28, 36] to downstream tasks. In this paper, we focus on transferring pre-trained models to video understanding tasks. ViFi-CLIP [29] demonstrates that direct fine-tuning exhibits good generalization capabilities across various settings. Open-VCLIP [38] constructs an open-vocabulary video model by interpolating the model weights and its optimization trajectory. Vita-CLIP [37] uses multi-level prompts to extract discriminative information. X-CLIP [24] introduces cross-frame attention and multi-frame integration modules for temporal modeling. OST [6] optimizes text knowledge by leveraging large language models to generate spatio-temporal descriptors, and proposes an optimal descriptor solver to enhance generalization. MoTE [44] inserts a mixture-of-temporal-experts into the visual encoder to capture diverse views of data bias. However, all of these methods require learning a large number of parameters and suffer from catastrophic forgetting of the original generalizable knowledge. Moreover, these methods are specifically designed for CLIP [28], making them unsuitable for the latest base model, i.e., ViCLIP [36]. ViCLIP is a pre-trained model based on CLIP, trained on 10 million video-text pairs. Our work addresses these challenges by designing an efficient tranfer leraning approach tailored for ViCLIP, demonstrating superior results across various settings.
3 Methods
3.1 Preliminaries
We adopt ViCLIP [36] as the pre-trained video-language foundation model in our method. ViCLIP is an improved version of CLIP that replaces the native attention in the visual encoder with spatiotemporal attention, while keeping other design components unchanged. It uses pre-trained CLIP weights for initialization. The model is trained on the InternVid-10M dataset [36], with the optimization objective being the contrastive loss between input video and text embeddings. ViCLIP is composed of two branches: a text branch with encoder $E_{t}$ and a vision branch with encoder $E_{v}$. These two branches enable it to understand and bridge the semantic gap between textual descriptions and visual content. In particular, a video $V\in\mathbb{R}^{T\times H\times W\times 3}$ will be fed into the encoder $E_{v}$ to obtain the video feature $x$ as follows:
$x_{0}$ $=\text{PatchEmbed}(V),$ (1)
$[c_{i},x_{i}]$ $=\mathcal{E}^{i}_{v}([c_{i-1},x_{i-1}]),\,i=1,2,...,L,$ (2)
$x$ $=\text{PatchProj}(c_{L}).$ (3)
Here, PatchEmbed first divides the input video $V$ into fixed-size patches, projecting them into feature embeddings. A learnable class token $c_{0}$ is then concatenated with these embeddings, forming $[c_{0},x_{0}]$, which is subsequently passed through $L$ transformer blocks, represented as $\{\mathcal{E}^{i}_{v}\}_{i=1}^{L}$. Finally, the class token $c_{L}$ from the last transformer block $\mathcal{E}^{L}_{v}$ is projected into the video feature $x$ via a projection layer PatchProj, positioning the feature in the shared vision-language space. In a similar manner, a text description $T$ is processed through the text encoder $E_{t}$ to obtain the text feature $w$ as follows:
$w_{0}$ $=\text{TextEmbed}(T),$ (4)
$w_{i}$ $=\mathcal{E}^{i}_{t}(w_{i-1}),\,i=1,2,...,L,$ (5)
$w$ $=\text{TextProj}(w_{L}).$ (6)
As shown, this process involves three steps: first, TextEmbed is used to tokenize and project the input text description into the initial word embedding $w_{0}$. Next, a series of transformer blocks $\{\mathcal{E}_{t}^{i}\}_{i=1}^{L}$ progressively abstracts the features, producing a refined embedding $w_{L}$ at the final layer. Finally, TextProj projects the output $w_{L}$ from the last transformer block $\mathcal{E}_{t}^{L}$ into the common vision-language space. Given these features, we compute the cosine similarity scores $sim(x,w)$ between the video and text descriptions across different domains or tasks to perform task-specific predictions.
Figure 2: The proposed Multi-modal Spatil-Temporal Adapter (MSTA) for the transformer-based multi-modal foundation models. Only the additional adapters are optimized, while the entire pretrained CLIP model remains frozen. To achieve a balance between the trade-off of discrimination and generalization, we selectively tune only a few higher layers ($\geq k$). MSTA utilizes shared weights across video and text representations, allowing the model to capture common cues from both branches.
3.2 Multi-modal Spatio-Temporal Adapter
Our work mainly focuses on generalization tasks (i.e., few-shot generalization [43, 19] and zero-shot generalization [34]), where the pre-trained multi-modal foundation models are first fine-tuned on a limited set of training samples, and then applied to recognize unseen instances. For such problems, a good instance representation should not only be discriminative but also exhibit strong generalization capabilities across different types of datasets. Typically, adding additional parameters helps better fit the training data distribution, thereby improving performance on seen tasks. However, models optimized for specific target distributions are often highly sensitive to external distribution shifts, which may lead to poor generalization performance in downstream tasks involving unknown video categories. While, in theory, expanding the training dataset to cover more unseen categories could mitigate this issue, the high computational and data acquisition costs associated with video data make this approach impractical. Therefore, it is crucial to develop a foundational model architecture with enhanced generalization capability.
Recently, parameter efficient fine tuning [17, 5] (PEFT) has proven to be an effective method for transfer learning in large models. Two representative approaches are LoRA [17] and AdaptFormer [5]. However, these methods are designed for uni-modal models and cannot be directly applied to multi-modal models. While it is possible to apply adapters separately to the two modalities, the lack of interaction between visual and textual information limits their performance. Therefore, we propose a new adapter-based efficient transfer learning framework as described below.
Holistic design. As shown in Fig. 2, different from most of existing methods that add adapters or tokens to the whole network or some lower layers, the new adapter A are partially added into a few higher-layers of both image and text encoders. Formally, for the image encoder V, we add our adapters $A_{v}$ from the $k$-th transformer block and modified Eq. (2) as follows:
As illustrated in Fig. 2, unlike most existing methods that introduce adapters or tokens across the entire network or into lower layers, our adapter $\mathcal{A}$ is selectively incorporated into only a few higher layers of both the video and text encoders. Specifically, for the video encoder $E_{v}$, we incorporate adapters $\{\mathcal{A}_{v}^{j}\}_{j=k}^{L}$ starting from the $k$-th transformer block, modifying Eq. (2) as follows:
$[c_{i},x_{i}]$ $=\mathcal{E}^{i}_{v}([c_{i-1},x_{i-1}]),\,i=1,2,\dots,k-1,$ (7)
$[c_{j},x_{j}]$ $=\mathcal{E}^{j}_{v}([c_{j-1},x_{j-1}])+\lambda\boxed{\mathcal{A}^{j}_{v}([c_{j-1},x_{j-1}])},$
$j$ $=k,k+1,\dots,L.$ (8)
Here, the portion inside the box represents the trainable blocks. The coefficient $\lambda$ is used to balance task-specific knowledge with general pre-trained knowledge. Notably, setting $\lambda=0$ reduces the model to the original transformer block, without incorporating any additional knowledge. Similarly, we add adapters $\{\mathcal{A}_{t}^{j}\}_{j=k}^{L}$ to the text encoder $E_{t}$ and modified Eq. (5) as follows:
$w_{i}$ $=\mathcal{E}^{i}_{t}(w_{i-1}),\quad i=1,2,\dots,k-1,$ (9)
$w_{j}$ $=\mathcal{E}^{j}_{t}(w_{j-1})+\lambda\boxed{\mathcal{A}^{j}_{t}(w_{j-1})},\,j=k,k+1,\dots,L.$ (10)
Figure 3: Overview of the proposed Multi-modal Spatil-Temporal Adapter (MSTA).
Specific design. According to  [16], the adapters in both the text branch and the video branch consist of down-projection layers (encoders), adapter layers, and up-projection layers (decoders). As shown in Fig. 3, to bridge the representations in both branches and reduce the semantic gap, we do not independently adapt the adapters in the video and text branches. Instead, we aggregate these bimodal signals through a shared projection layer. During fine-tuning, this shared feature space receives gradients from both modalities, thereby optimizing their alignment. Additionally, for the video branch, we designed two types of up-projection layers: one spatial up-projection layer and one temporal up-projection layer, to further enhance the adaptability to spatiotemporal features. Formally, this process can be summarized as follows:
$\mathcal{A}^{k}_{v}(h_{k})$ $=\mathbf{W}_{v}^{ku-s}\cdot h_{k}+\mathbf{W}_{v}^{ku-t}\cdot h_{k},$ (11)
$h_{k}$ $=\sigma(\mathbf{W}^{ks}\cdot\sigma(\mathbf{W}_{v}^{kd}\cdot[c_{k},x_{k}])).$ (12)
A similar process is added to text encoder as follows:
$\mathcal{A}^{k}_{t}(w_{k})=\mathbf{W}_{t}^{ku}\cdot\sigma(\mathbf{W}^{ks}\cdot\sigma(\mathbf{W}_{t}^{kd}\cdot w_{k})).$ (13)
Here, $\mathbf{W}^{ku}$ and $\mathbf{W}^{kd}$ represent the ”Up” and ”Down” projection layers of the $k$-th layer, as illustrated in Fig. 3, with the modality branch indicated by the subscripts. $\mathbf{W}_{v}^{ku-s}$ is the spatial up-projection layer implemented by a linear layer, while $\mathbf{W}_{v}^{ku-t}$ is the temporal up-projection layer implemented by a 3D convolution layer. $\mathbf{W}^{ks}$ denotes the $k$-th projection layer, which is shared across different branches.
3.3 Spatio-Temporal Description-guided Consistency Constraint
[…]
$\mathcal{L}_{\rm CC}=2-\frac{w^{c}\cdot D^{c}_{s}}{||w^{c}\cdot D^{c}_{t}||||D^{c}_{s}||}-\frac{w^{c}}{||w^{c}||||D^{c}_{t}||},$ (14)
where $w^{c}$ indicates the text embedding for class $c$ of the pre-trained branch, and $D^{c}_{s}$ represents the mean embedding of the spatio descriptions for class $c$, obtained by averaging the features of the input descriptions processed through the pre-trained branch. Similarly, $D^{c}_{t}$ is computed for the temporal descriptions. This consistency constraint loss is combined with a supervised loss to create the final loss function. The supervised loss is given by:
$\mathcal{L}_{\rm CE}=-log\frac{{\rm exp}(sim(x,w^{c})/\tau}{\sum_{k=1}^{C}{\rm exp}(sim(x,w^{k})/\tau},$ (15)
where $\tau$ is a temperature parameter controlling the sharpness of similarity scores. By combining these losses with a weighting factor $\alpha$, the final loss function for MSTA becomes:
$\mathcal{L}=\mathcal{L}_{\rm CE}+\alpha\mathcal{L}_{\rm CC}.$ (16)
Table 1: Implementation details of MSTA
Dataset Batch size Learning rate Training epochs $\alpha$ $\lambda$ Dims $N$ Dropout Layers
Base-to-novel generalization
K-400 32 1.e-03 11 1.0 0.005 256 2 0.1 1-12
SSv2 32 1.e-03 11 1.0 0.005 256 2 0.1 1-12
HMDB-51 32 1.e-03 11 1.0 0.005 256 2 0.1 1-12
UCF-101 32 1.e-03 11 1.0 0.005 256 2 0.1 1-12
Few-shot learning
SSv2 32 1.e-03 50 1.0 0.005 128 2 0.1 8-12
HMDB-51 32 1.e-03 50 1.0 0.005 128 2 0.1 8-12
UCF-101 32 1.e-03 50 1.0 0.005 128 2 0.1 8-12
Zero-shot Transfer
K-400 256 8.e-03 100 1.0 0.001 128 2 0.1 1-12
Few-shot learning
K-400 256 8.e-03 100 1.0 0.001 128 2 0.1 1-12
Table 2: Base-to-novel generalization: We compare the generalization ability of MSTA with models that adapt ViCLIP for video tasks. “HM” is the harmonic mean of base and novel accuracy, providing the trade-off between adaption and generalization. ”Zero-ViCLIP” refers to zero-shot ViCLIP, which directly uses the pre-trained ViCLIP for inference.
K-400 HMDB-51 UCF-101 SSv2
Method Base Novel HM Base Novel HM Base Novel HM Base Novel HM
Zero-shot pre-trained models
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 55.1 55.2 55.1 50.6 48.1 49.3 78.0 63.8 70.2 4.8 5.4 5.1
Adapting pre-trained models
ActionCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(arXiv^{\prime}22)}}$ 61.5 47.2 53.4 69.3 38.3 49.3 90.5 58.7 71.2 13.5 10.4 11.7
XCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 74.7 56.7 64.5 69.9 45.3 55.0 90.7 59.5 71.9 8.9 6.8 7.7
A5${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 69.9 38.4 49.6 46.9 16.5 24.4 90.9 40.8 56.3 8.6 6.4 7.3
Tuning pre-trained models
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 76.4 63.0 69.1 71.5 55.3 62.4 92.3 67.9 78.2 13.2 10.8 11.9
ViFi-CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}23)}}$ 76.8 61.3 68.2 74.5 53.9 62.5 92.5 68.5 78.7 16.5 12.4 14.2
OST${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}24)}}$ 75.4 59.6 66.6 75.1 34.4 47.2 96.1 68.9 80.3 15.3 11.5 13.1
ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 78.5 65.3 71.3 77.9 52.4 62.7 95.1 71.5 81.6 19.4 15.2 17.0
Efficient transfer pre-trained models
Zero-ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 68.7 63.4 65.9 62.5 48.9 54.9 81.7 69.8 75.3 9.3 8.9 9.1
+ AdaptFormer${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 77.7 65.6 71.1 77.5 54.9 64.3 94.9 72.7 82.3 20.4 14.6 17.0
+ LoRA${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 77.6 65.3 70.9 77.2 54.6 64.0 95.1 72.3 82.1 19.5 13.6 16.0
+ MSTA 78.2 66.2 71.7 77.2 57.5 65.9 95.6 72.8 82.7 21.9 16.3 18.7
+ MSTA + $\mathcal{L}_{\rm CC}$ 78.5 66.5 72.0 77.5 57.9 66.3 96.0 72.9 82.9 22.2 16.5 18.9
Table 3: Comparisons with state-of-the-art methods for few-shot video recognition on HMDB51, UCF101 and Something-Something V2. We scaled up the task to categorize all categories in the dataset with only a few samples per category for training. Here $K$ denotes training samples for each class. We report Top-1 accuracy using multi-view inference.
Method HMDB-51 UCF-101 SSv2
$K=2$ $K=4$ $K=8$ $K=16$ $K=2$ $K=4$ $K=8$ $K=16$ $K=2$ $K=4$ $K=8$ $K=16$
Zero-shot pre-trained models
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 37.2 37.2 37.2 37.2 62.8 62.8 62.8 62.8 2.8 2.8 2.8 2.8
Adapting pre-trained models
XCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 53.5 57.8 62.9 64.5 49.3 76.2 84.1 91.6 4.4 5.2 6.6 11.1
A5${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 41.2 51.3 56.5 62.7 71.7 80.3 85.9 90.3 4.5 5.3 6.3 9.9
Tuning pre-trained models
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 57.8 61.2 65.6 66.7 81.8 86.3 90.1 92.5 6.7 7.3 9.0 12.2
ViFi-CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}23)}}$ 57.5 62.9 64.7 66.6 80.5 85.3 90.2 93.2 6.5 7.7 8.3 12.9
OST${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}24)}}$ 60.0 62.5 65.6 67.3 83.0 88.4 91.3 93.9 7.3 8.4 8.5 11.5
MoTE${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}24)}}$ 61.0 63.7 66.9 68.5 88.1 90.7 92.2 93.7 7.4 8.7 9.8 12.5
ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 53.7 60.4 64.5 70.3 83.0 88.0 92.1 93.2 8.7 9.7 11.6 15.4
Efficient transfer pre-trained models
Zero-ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 47.8 47.8 47.8 47.8 71.0 71.0 71.0 71.0 5.1 5.1 5.1 5.1
+ AdaptFormer${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 60.0 61.8 66.6 70.1 84.8 88.7 91.7 93.3 8.5 10.1 11.5 15.2
+ LoRA${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 59.4 61.4 66.5 69.4 83.5 88.3 90.9 92.7 8.4 9.5 11.1 14.9
+ MSTA 60.1 62.2 66.8 69.9 85.1 89.0 92.3 93.5 9.0 10.1 12.7 16.6
+ MSTA + $\mathcal{L}_{\rm CC}$ 61.2 62.3 67.1 70.4 86.1 90.2 92.7 94.8 9.1 10.4 13.4 17.5
4 Experiments
4.1 Datasets
Kinetics-400 [1] is a large-scale video dataset containing approximately 240,000 training videos and 20,000 validation videos across 400 human action categories, with an average video length of 10 seconds. Due to its high quality, it has become one of the most popular benchmarks for video recognition.
Kinetics-600 [2] is an extension of Kinetics-400, with approximately 392,000 training videos, 30,000 validation videos, and 60,000 test videos in 600 human action categories. It introduces 220 new action categories not present in Kinetics-400. We evaluate the zero-shot performance on these 220 new categories and use three splits provided by previous work [4]. The test set is used for evaluation, and we report the average performance across the three splits.
UCF-101 [31]is an action recognition dataset containing 13,320 videos in 101 action categories, collected from YouTube. It has three official splits for training and validation data.
HMDB-51 [20] contains 7,000 videos across 51 action categories, collected from movie clips and web videos. The dataset has three official splits, with 3,570 training videos and 1,530 validation videos in each split.
Something-Something V2 [13] is a temporally demanding dataset that requires fine-grained temporal understanding. It includes 220,000 videos across 174 action categories.
4.2 Experimental Setup
Architecture. We employ the ViCLIP [36] pre-trained ViT-B/16 in our experiments. For the zero-shot transfer, few-shot learning, base-to-novel generalization and fully-supervised settings, we introduced the proposed multi-modal spatio-temporal adapter starting from the $k=1/8/1/1$-th transformer block to the last one in both the language and vision branches, respectively.
Implementation Details. Table 1 outlines the implementation specifics for efficient transfer video recognition. All modules in MSTA are initialized using kaiming initialization [14]. We use AdamW [22] as the optimizer, with a weight decay of 0.001 and Adam’s $\beta_{1}$, $\beta_{2}$ set to 0.9 and 0.98, respectively. The linear warm-up consists of 5 epochs. For data augmentation during training, we apply ColorJitter (P=0.8), GrayScale (P=0.2), RandomResizedCrop, and FLIP (flip ratio=0.5). All experiments are conducted on 8 Nvidia Tesla-A100-80G GPUs.
Evaluation Protocols. (1) Zero-shot: Following previous works [24, 6], we evaluate zero-shot performance on UCF-101 [31], HMDB-51 [20], Kinetics-600 [2]. In zero-shot setting, we test using single view with 8 frames. (2) Few-shot and base-to-novel generalization: We evaluate on Kinetics-400 [1], UCF-101 [31], HMDB-51 [20], and SomethingSomething V2 [13]. Following [32], we adopt a multi-view for evaluation. (3) Fully-supervised: We evaluate the fully-supervised performance on Kinetics-400 [1], using 4 clips with 3 crops (i.e. 4$\times$3 views) per video [24]. Each view contains 8 sparsely sampled frames.
Baseline models. We compare our approach to efficient transfer learning methods [4, 27, 24, 18, 37, 29, 44, 6] that based on CLIP for video tasks. Since these methods are specifically designed for CLIP, they cannot be directly applied to ViCLIP. Therefore, we employed common parameter-efficient fine-tuning methods [16, 17, 5] on ViCLIP to enable a more comprehensive comparison.
4.3 Main Results
Base-to-novel generalization. In Table 2, we compare our method with the state-of-the-art results under the base-to-novel setting. Our method achieves new state-of-the-art results on all datasets. Compared to OST [6] and ActionCLIP [34], which model video-specific inductive biases through full fine-tuning, our approach provides better base accuracy with minimal design modifications and demonstrates a significant improvement in novelty accuracy. Our method achieves a better trade-off between base and novel accuracies across all datasets and obtains the overall best harmonic mean. Moreover, it shows enhanced capability in understanding scene dynamics, particularly on temporally challenging datasets like SSv2. Specifically, compared to OST, we improve the recognition accuracy of the novel category on SSv2 from 11.5 to 15.8 (an increase of 30 37%). Our method also outperforms classic PEFT approaches (i.e., AdaptFormer [5], LoRA [17]), validating the effectiveness of MSTA and consistency constraints in enhancing generalization capabilities.
Table 4: Comparisons with state-of-the-art methods for zero-shot video recognition on HMDB51, UCF101 and Kinetics-600. We report Top-1 and Top-5 accuracy using single-view inference.
Method Frames HMDB-51 UCF-101 K600
Zero-shot pre-trained models
ER-ZSAR${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICCV^{\prime}21)}}$ 16 35.3 $\pm$ 4.6 51.8 $\pm$ 2.9 42.1 $\pm$ 1.4
JigsawNet${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 16 38.7 $\pm$ 3.7 56.0 $\pm$ 3.1 -
Adapting pre-trained models
Vanilla CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICML^{\prime}21)}}$ 32 40.8 $\pm$ 0.3 63.2 $\pm$ 0.2 59.8 $\pm$ 0.3
XCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 32 44.6 $\pm$ 5.2 72.0 $\pm$ 2.3 65.2 $\pm$ 0.4
A5${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 8 / 32 44.3 $\pm$ 2.2 69.3 $\pm$ 4.2 55.8 $\pm$ 0.7
Vita-CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}23)}}$ 32 48.6 $\pm$ 0.6 75.0 $\pm$ 0.6 67.4 $\pm$ 0.5
Tuning pre-trained models
ViFi-CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}23)}}$ 32 51.3 $\pm$ 0.7 76.8 $\pm$ 0.8 71.2 $\pm$ 1.0
OST${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}24)}}$ 8 54.9 $\pm$ 1.1 77.9 $\pm$ 1.3 73.9 $\pm$ 0.8
ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 8 51.9 $\pm$ 0.7 77.1 $\pm$ 0.9 71.5 $\pm$ 1.1
Efficient transfer pre-trained models
Zero-ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 8 39.8 $\pm$ 0.9 62.6 $\pm$ 0.5 59.2 $\pm$ 0.4
+ AdaptFormer${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 8 54.1 $\pm$ 0.9 77.2 $\pm$ 0.6 72.8 $\pm$ 0.8
+ LoRA${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 8 54.2 $\pm$ 0.5 77.3 $\pm$ 0.8 73.8 $\pm$ 0.7
+MSTA 8 55.3 $\pm$ 0.7 77.8 $\pm$ 0.9 74.1 $\pm$ 0.7
+MSTA+$\mathcal{L}_{\rm CC}$ 8 55.8 $\pm$ 0.8 78.7 $\pm$ 0.9 74.5 $\pm$ 0.5
Few-shot learning. In Table 3, we conducted a challenging full-shot few-shot video recognition task, which requires rapid adaptation to new categories with limited samples, demanding both specialization and generalization. Our approach achieved 9 out of 12 best performances across different sample settings for all datasets, demonstrating strong learning capability and transferability. Notably, compared to the second-best method, MoTE [44], we used only 10% of its parameter count (i.e., 8.7M vs. 88M).
Zero-shot transfer. We present our zero-shot video recognition results in Table 4 and compare our approach with current state-of-the-art methods. First, the model is fine-tuned on the Kinetics-400 dataset [1] and directly evaluated on downstream datasets to assess its generalization capability to unseen classes. As shown in the table, our method significantly outperforms the conventional uni-modal zero-shot video recognition pipeline [4, 27]. Furthermore, we compare our approach with methods [24, 18, 37, 6] that adapt the CLIP model for zero-shot recognition using Kinetics-400, as well as methods that fine-tune ViCLIP with classic PEFT approaches (e.g., AdaptFormer [5], LoRA [17]). We observe consistent improvements across all datasets compared to these methods.
Fully-supervised learning. We also conducted fully supervised experiments on the large-scale video benchmark dataset Kinetics-400 to validate the effectiveness of our method in supervised settings. As shown in Table 5, our method achieved the best performance. Specifically, our method outperforms Vita-CLIP [37] by approximately 2% on K400.
Table 5: Comparisons with state-of-the-art methods for fully-supervised video recognition on Kinetics-400. We report Top-1 and Top-5 accuracy using single-view inference.
Method Frames Tunable K-400
Param Top-1 Top-5
XCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ECCV^{\prime}22)}}$ 8 132 82.3 95.8
Vita-CLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}23)}}$ 8 39 80.5 95.9
OST${}_{{\color[rgb]{0.5,0.5,0.5}\rm(CVPR^{\prime}24)}}$ 8 149.6 82.0 95.8
ViCLIP${}_{{\color[rgb]{0.5,0.5,0.5}\rm(ICLR^{\prime}23)}}$ 8 124.3 79.9 95.1
+ AdaptFormer${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 8 7.9 80.2 95.5
+ LoRA${}_{{\color[rgb]{0.5,0.5,0.5}\rm(NeurIPS^{\prime}22)}}$ 8 9.4 80.3 95.4
+ MSTA 8 8.7 81.6 95.8
+ MSTA+$\mathcal{L}_{\rm CC}$ 8 8.7 82.2 96.2
Table 6: Ablation studies over 4 datasets used in base-to-novel generalization setting.
Model Variants Base Novel HM
Only L-Adapter 66.1 51.5 57.9
Only V-Adapter 65.7 51.7 57.9
w/o Shared Layers 68.0 52.9 59.5
Full 68.6 53.5 60.1
(a) Performance with Different Model Variants
Dims Base Novel HM
64 67.5 52.2 58.9
128 68.2 53.1 59.7
256 68.6 53.5 60.1
512 68.7 53.3 60.0
(b) Dimensions of Shared Layers
$\lambda$ Base Novel HM
0.001 67.7 53.1 59.5
0.005 68.6 53.5 60.1
0.01 68.5 52.9 59.7
0.05 67.1 40.3 50.4
(c) Scaling Factor
$\alpha$ Base Novel HM
0.1 68.1 53.2 59.7
1.0 68.6 53.5 60.1
5.0 68.5 53.0 59.8
(d) Weighting Factor
$N$ Base Novel HM
2 68.6 53.5 60.1
4 68.8 53.0 59.9
8 68.6 52.9 59.7
(e) Discription Number
Layer 1$\rightarrow$6 1$\rightarrow$12 7$\rightarrow$12 8$\rightarrow$12 10$\rightarrow$12
Base 67.4 68.6 68.1 68.7 65.9
Novel 52.9 53.5 53.0 53.1 51.9
HM 59.3 60.1 59.6 59.9 58.1
(f) MSTA inserted layers.
4.4 Ablation Studies
We conduct ablation studies on base-to-novel settings in Table 6 to investigate the learning capacity and generalizability of our model in different instantiations.
Component-wise analysis of MSTA. We began by examining the impact of various adapter configurations. Specifically, we compared the performance of uni-modal adapters, which are adapters applied exclusively to either the vision (V-) or language (L-) branch, against the performance of our full MSTA model, which incorporates adapters in both modalities. Our findings, as presented in Table 4a, clearly indicate that the uni-modal adapters are outperformed by the full MSTA model. This suggests that the integration of both visual and textual information is crucial for optimal performance. Moreover, we discovered that omitting the shared projection layers (w/o Shared Layers) results in a noticeable performance drop. When these shared layers are included, we observe an increase in the harmonic mean (HM) from 59.5 to 60.1 across the four datasets.
Dimension of the Shared Layer. The dimension of the shared layer in our MSTA determines the number of parameters required to extract relationships from the features of the two modalities. We conducted an ablation study by systematically varying the dimension of the shared layer to assess its impact. As shown in Table 6(b), accuracy for base categories peaks as the intermediate dimension increases, while accuracy for novel categories saturates around 256. This may be due to larger dimensions introducing more trainable parameters, which increases the risk of over-fitting.
Scaling Factor $\lambda$. The scaling factor $\lambda$ is a critical hyperparameter that controls the balance between general and task-specific features within the MSTA. As shown in Table 6(c), we thoroughly evaluated various scaling factors. Our results reveal that a scaling factor of $\lambda=0.005$ provides the best performance trade-off (HM) between base and novel categories. This optimal balance is essential for enabling the model to quickly adapt to base categories without sacrificing its ability to generalize to novel categories. In contrast, larger scaling factors, while improving performance on base categories, can hinder the model’s performance on novel categories. Smaller scaling factors, on the other hand, can limit the model’s adaptability to downstream tasks.
Weighting Factor $\alpha$. The weighting factor $\alpha$ associated with the consistency constraint loss is a key hyperparameter that influences model performance. As shown in Table 6(d), the choice of $\alpha$ has a significant impact on accuracy. The model achieves peak performance when $\alpha=1.0$. Deviations from this value, whether by increasing or decreasing $\alpha$, lead to a decline in performance. This finding underscores the importance of carefully tuning the consistency constraint to maintain a balance between adaptability and generalization capabilities.
Variants of Adding MSTA. We evaluate the impact of adding MSTA to different encoder layers, with results summarized in Table 6(f). The performance of MSTA improves consistently as the number of added layers increases. Furthermore, given the same number of layers, MSTA performs better when placed higher in the network (i.e., farther from the input). For example, inserting MSTA into layers 7 to 12 yields 0.3 higher HM than inserting into layers 1 to 6, despite both configurations using six MSTA layers. Notably, the insertion of MSTA in layers 8 to 12 delivers the best performance for base class tasks, which is particularly advantageous for few-shot learning scenarios where rapid adaptation to new categories is essential.
Impact of different descriptions. We investigate the impact of spatiotemporal descriptions on the performance of our proposed method. The results presented in Table 7 show that each description complements the others, highlighting the importance of both spatiotemporal descriptions for recognition tasks. Additionally, we observe that the temporal descriptions have a more pronounced effect on performance compared to the spatiotemporal descriptions.
Table 7: Study on different descriptions.
Description Base Novel H
w/o DES 68.2 53.2 59.77
w/o Spatio 68.4 53.4 59.97
w/o Temporal 68.8 53.3 60.06
Full 68.6 53.5 60.12
Cost Analysis. We analyze the additional costs of our method during training and inference in Table 8. Latency is measured in our baseline training setup, while throughput is evaluated using the largest batch size that can be processed without exceeding the memory limit of a single NVIDIA A100-80G. Our pipeline incurs only an additional 0.1× training time and results in a throughput reduction of approximately 5%, which is acceptable considering the performance gains achieved.
Table 8: Additional cost analysis of our method, we report step latency during training, and throughput (TP) during inference. We refer to Top-1 as zero-shot accuracy on Kinetics-600.
Description Top-1 (%) Latency (s) TP (video/s)
ViCLIP [36] 71.5 0.45 (1.0x) 69.3(1.0x)
Ours 74.5 0.51 (1.1x) 65.7 (0.95x)
5 Conclusion
In conclusion, this paper presents a novel Multi-modal Spatio-Temporal Adapter (MSTA) for efficient transfer learning in video-language foundation models. The MSTA is designed to enhance the alignment between text and video representations, thereby achieving a balance between general knowledge and task-specific knowledge. Our method introduces a spatio-temporal description-guided consistency constraint to mitigate over-fitting and enhance generalizability, which is particularly beneficial in data limited scenarios. Through extensive experiments across various tasks, our MSTA demonstrates outstanding performance and achieves state-of-the-art results while using a minimal number of trainable parameters. The comprehensive ablation studies further validate the effectiveness and scalability of our proposed method. Overall, our work contributes significantly to the field of video action recognition by providing an efficient and effective transfer learning approach that preserves the generalizability of pre-trained models while adapting them to specific downstream tasks.
6 Limitation and Broader Impact
Limitation. Although our method achieves excellent results across various video recognition settings, there are still areas for improvement. For example, MSTA connects the gradients of visual and textual features. Despite using gradient scaling to prevent instability, we believe that the backpropagation performed under this setup is not optimal. Therefore, exploring better ways to train this type of module could further enhance MSTA’s performance. On the other hand, using spatiotemporal descriptions generated by LLMs for constraints may lead to performance degradation due to the lower quality of these descriptions. In future work, we will further investigate how to leverage LLMs to improve the efficient transfer of video-language models, aiming to explore the generality and versatility of our approach.
Broader Impact. The adaptability of foundational models to various downstream tasks has become a major focus in the current field of machine learning [17, 5, 25]. We believe that retaining the existing capabilities of foundational models while integrating new knowledge is an important issue worth exploring in depth. We hope this research provides new insights and approaches for the broader and long-term application of foundational models. Our work focuses on video recognition tasks, which have widespread applications in real-world scenarios such as surveillance. However, before these technologies can be deployed, it is crucial to carefully address potential issues related to privacy protection and rights.
7 Pseudo-code
The pseudo-code implementation of MSTA is provided below, along with a concise definition of MSTA and its insertion method.
Algorithm 1 Implementation of Text Adapter & Shared Adapter in PyTorch-like style
  def _build_adapter(d_model, n_layers, l_start, l_end, mid_dim, dropout_rate=0.1):
        adapter = [None] * (n_layers + 1)
        for i in range(l_start, l_end+1):
            if mid_dim == d_model:
                adapter[i] = nn.Sequential(
                    nn.Linear(d_model, mid_dim),
                    nn.ReLU()
            else:
                adapter[i] = nn.Sequential(OrderedDict([
                    (”down”, nn.Sequential(
                    nn.Linear(d_model, mid_dim),
                    nn.ReLU()
                    (”dropout”, nn.Dropout(p=dropout_rate)),
                    (”up”, nn.Linear(mid_dim, d_model))
        adapter = nn.ModuleList([a for a in adapter])
        return adapter
Algorithm 2 Implementation of Visual Adapter in PyTorch-like style
  def _build_visual_adapter(d_model, n_layers, l_start, l_end, mid_dim, dropout_rate=0.1):
        adapter = [None] * (n_layers + 1)
            for i in range(l_start, l_end+1):
                adapter[i] = nn.Sequential(OrderedDict([
                (”down”, nn.Sequential(
                nn.Linear(d_model, mid_dim),
                nn.ReLU()
                (”dropout”, nn.Dropout(p=dropout_rate)),
                (”up”, nn.Linear(mid_dim, d_model)),
                (”temporal_up”, nn.Conv3d(mid_dim, d_model, kernel_size=(3,1,1), stride=(1, 1, 1),
                padding=”same”, groups=mid_dim, dilation=(1, 1, 1))),
        adapter = nn.ModuleList([a for a in adapter])
        return adapter